# Practical solutions to the relative pose of three calibrated cameras

### v4

@@ 5 Conclusion
We consider the highly challenging problem of relative pose estimation of three calibrated cameras from four correspondences. We propose novel approaches that solve the problem by utilizing solvers based on approximate geometry. The best performing solver uses a simple, yet novel strategy by using mean coordinates of three input points and points in their vicinity as an approximate correspondences. Extensive experiments show that our solvers achieve state-of-the-art performance on a large variety of real scenes. At the same time, our solvers are simple to implement, especially compared to the current state-of-the-art [22].
Acknowledgements. This work was funded by the Czech Science Foundation (GAČR) JUNIOR STAR Grant No. 22-23183M (supporting C.T., Y.D., and Z.K.), the Grant Agency of the Czech Technical University in Prague grant no. SGS23/173/OHK3/3T/13 (supporting C.T.), the EU NextGenerationEU through the Recovery and Resilience Plan for Slovakia under the project ”InnovAIte Slovakia, Illuminating Pathways for AI-Driven Breakthroughs" No. 09I02-03-V01-00029, the TERAIS project – a Horizon-Widera-2021 program of the European Union under the Grant agreement number 101079338 (supporting V.K. and Z.B.H.), and the EU Horizon 2020 project RICAIP grant agreement No. 857306 (supporting T.S.). Part of the research results was obtained using the computational resources procured in the national project National competence centre for high performance computing project code: 311070AKF2, funded by European Regional Development Fund, EU Structural Funds Informatization of society, Operational Program Integrated Infrastructure.
@@
This supplementary material provides additional details and experimental results promised in the main paper: Sec. 6 provides the proof on the bound of the epipolar error of the mean point correspondence mentioned in Sec. 3.1 of the main paper, additional synthetic and noise experiments that were discussed in Sec. 4 of the main paper, accuracy-speed trade-off results for two-view approximate geometry inside RANSAC (see Sec. 4 of the main paper), and additional details and plots on more scenes for the mean point correspondence accuracy (see Sec. 4 of the main paper). Sec. 7 contains ablation studies to validate our choices regarding the modifications discussed in Sec. 3.2 of the main paper. Sec. 8 provides results for the accuracy-speed trade-off for individual scenes from the PhotoTourism [23] and Cambridge Landmarks [24] datasets, evaluations of the three-view solvers for different thresholds inside RANSAC (see Sec. 4 of the main paper), runtimes of the proposed and state-of-the-art solvers for the 4p3v problem, semi-synthetic experiments for increasing outliers ratios on a scene from PhotoTourism [23], details and results on an alternative evaluation measure (see Sec. 4 of the main paper), and results with GC-RANSAC (see Sec. 4 of the main paper). 44footnotetext: ∗ Equal contribution
@@ 6 Approximate camera geometry
In this section, in addition to the experiments presented in Sec. 4 of the main paper (paragraph “Approximate camera geometry"), we present additional experiments and results to support our idea of estimating approximate geometry in the first two views. We start with the proof on the bound of the epipolar error of the mean-point correspondence used in the proposed 4p3v(M)-based solvers (see Sec 3.1 of the main paper). Next, Sec. 6.2 discusses why the mean-point correspondence provides an additional constraint (compared to the original point correspondences) that can be used to estimate the essential matrix. Then, to further assess the accuracy of the two-view variants of our approximate solvers (outside of RANSAC), i.e., 4p(A), 4p(M), and 4p(M), we design two additional synthetic experiments (see Sec. 6.3). The goal of these synthetic experiments is to study how the accuracy of approximate solutions varies with varying properties of the scene and the cameras. Moreover, Sec. 6.3 provides a synthetic noise experiment on data extracted from a real scene from the ETH3D dataset [51], outside of RANSAC for the three-view solvers. Lastly, Sec. 6.4 contains a speed-accuracy evaluation of the two-view variants of the proposed solvers, inside Poselib RANSAC, and in Sec. 6.5 we study the accuracy of the mean point correspondence (as in Fig. 3 of the main paper).
@@ 6.1 Proof of the bounds on the epipolar error
While the mean point correspondence used in the 4p3v(M)-based solvers can provide a good approximation of a correct correspondence, such a correspondence can be noisy. Note that all state-of-the-art 4p3v solvers (including 4p3v(HC) [22] and the solver from [37]) rely on certain approximations without establishing theoretical proofs to quantify their accuracy. In the 4p3v(HC) solver [22], the failures that appear quite often are usually the results of tracking a geometrically incorrect solution inside the homotopy continuation method.555The solver is tracking only one from 272 solutions of the relaxed version of the 4p3v problem and this solution does not need to be a geometrically correct one. Thus, this solution can be arbitrarily far from the correct solution. The solver from [37] requires sampling epipoles from a -degree curve on which the true epipole must lie. For any selected point on the curve of epipoles, the error of the sampled epipole is not bounded, since the true epipole can lie anywhere on the unbounded curve. The curve is unbounded since the epipole can be located arbitrarily far away from the image center based on the relative pose of the two cameras (up to infinity for sideways motion).
In contrast, the error of our mean point correspondence is bounded. The correspondence can be seen as a correspondence of points that are projections of the mean point of three 3D points. In this case, the error of both projections and can be computed as an error that is introduced by approximating the perspective projection using the para-perspective projection. The approximation error introduced by the para-perspective projection is studied in the literature and can be found e.g. in [59, 21].
@@ 6.2 Mean-point constraint
As already mentioned in the main paper, under the assumption of a para-perspective projection, i.e., of affine geometry, the mean point of three 3D points is projected to the mean points of the 3D points’ projections in both images [59]. Thus, the mean point correspondence does not add a new constraint if used to estimate an affine camera. This can be easily shown. In the case of affine cameras, the essential matrix has the form
@@ 6.3 Synthetic Experiments
The error of the relative poses estimated with the proposed approximate 4p3v(M)-based and 4p3v(A)-based solvers depends on many aspects, e.g., the baseline and the view angles of the cameras w.r.t. the three points used to compute the mean point correspondence, the depth of these points, the size and shape of the triangles defined by the three points, the type of motion of the cameras, the depth of the scene and the distance of cameras from the scene, the level of noise in the correspondences, etc. Isolating the impact of the individual aspects, e.g., through experiments on synthetic data, is highly non-trivial (e.g., how to generate realistic synthetic scenarios that allow conclusions to generalize to real-world scenarios) and analysing the co-dependencies between different aspects on the overall performance seems to need a paper on its own. Moreover, the effect of approximation introduced by using para-perspective projection was already studied in the literature [59, 21].
@@
Additionally, we test the performance (outside of RANSAC) of our proposed approximate solvers and the state-of-the-art solvers for the 4p3v problem w.r.t. increasing image noise added to ground-truth correspondences extracted from a scene from the ETH3D dataset [51]. As an evaluation metric for the two-view geometry, we use the pose error measured as [23].
@@
Increasing distance of cameras to the 3D scene. It is known that the quality of the affine approximation, i.e., the approximation of the perspective projection using the para-perspective projection, depends on the distance of the points from the camera [59]. Thus in the first experiment, we evaluate the performance of all solvers w.r.t. increasing distance of the cameras to the 3D scene.
@@
Noise experiments. We test the performance of our solvers and the state-of-the-art solvers w.r.t. increasing image noise. We used the SfM model of the botanical garden scene (randomly selected from all scenes) from the ETH3D dataset [51] to obtain instances of 5 points in three views by identifying images in the scene that share 3D points. Perfect noise-free image correspondences are generated by projecting the 3D points into the images. We then add increasing amounts of normally distributed noise to these correspondences. We generated more than 1k instances. Note that the 4p3v(HC) solver was trained on the ETH3D dataset [51].
@@
The results for increasing noise in the image points are shown in Fig. 9. The figure shows boxplots of the pose errors measured in the same way as in our experiments in the main paper (cf. Sec. 4 in the main paper), i.e., as .666Here is the error of the estimated relative rotation between cameras and , computed as the angle in the axis-angle representation of , and is the error of the estimated translation computed as the angle between the two unit vectors corresponding to the translations [23]. The errors are zoomed into an interesting interval and are shown as functions of varying Gaussian noise from px to px.
@@
Due to the approximations used in our proposed 4p3v(M)-based and 4p3v(A)-based solvers, these solvers exhibit non-zero errors for zero noise. However, at noise levels px, our 4p3v(M)+R solvers return comparable or even better (w/ ENM) results than the 5pt+P3P solver. For noise px, also the 4p3v(M)+R solver with ENM returns slightly more accurate poses than the 5pt+P3P solver with ENM. In general, the effect of increasing image noise is less visible for approximate 4p3v(M)-based and 4p3v(A)-based solvers. In this case, the error of the approximation is dominating the error introduced by the noise in the image correspondences. While for 4p3v(A)-based solvers the approximation error is dominant at all noise levels, for 4p3v(M)-based solvers, at noise px, the error introduced by the approximate mean point correspondence (and points in their vicinity in -based solvers) is suppressed by the error introduced by noise in the remaining point correspondences.777This can be seen from the comparable pose accuracy of the 5pt+P3P and 4p3v(M) solvers for px noise, and the comparable accuracy of the 5pt+P3P and 4p3v(M) solvers for px noise. Note that, although the pose errors for the 4p3V(A)+R+ENM solver are higher than those of the rest of the solvers, as shown in our real experiments, this solver still returns reasonably low errors to provide local optimization (LO) within RANSAC with a good initialization in real-world settings. Further, note that the 5pt+P3P solver samples one more point (real correspondence) in the first two cameras, and these points are affected only by the considered noise. This shows that the mean point correspondence used in the 4p3v(M)-based solver is a good approximation to a real correspondence. The recent state-of-the-art 4p3v(HC) solver [22] is failing in about 50% of the instances for noiseless data, even though the solver was trained on the ETH3D dataset. Thus, the median errors are significantly larger than the median errors of the remaining solvers.
@@ 6.4 Two-view approximate solutions inside RANSAC
Fig. 10 shows the speed-accuracy evaluation of the solvers for the problem of two view relative pose estimation evaluated using PoseLib RANSAC [28] on 12 scenes from the Phototourism dataset [23] (excluding St. Peter’s Square) with pairwise point correspondences obtained using [10, 33]. The statistic reported is the AUC@10∘ of the pose error for a varied number of RANSAC iterations () and a fixed epipolar threshold of 5px. The upper figure is a zoom-in of the lower figure to an interesting interval, where differences between the solvers are more visible. Although all proposed solvers (except the 4p(A) solver without ENM) have a performance comparable to that of the state-of-the-art two view 5pt solver, the 5pt solver is the best performing one for the two-view scenario. This result is not surprising, given the well-known good performance (in terms of speed and accuracy) of the 5pt solver and not very high outlier contamination of the data (for which sampling one point less would have potentially had a more visible effect). It also indicates that the proposed modifications, i.e., the filtering +F the and refinement +R, for the three-view scenario are important and are making the proposed 4p3v approximate solvers practical and more precise than the 5pt+P3P sovler.
@@ 6.5 Accuracy of the mean point correspondence
Fig. 3 in the main paper showed results obtained by establishing correspondences between the mean of the triangle in one image and various points in the triangle in the second image. We expressed points in the second triangle via their barycentric coordinates and uniformly sample barycentric coordinates , such that (ensuring points inside the triangle). The 3rd coordinate is given as . For each correspondence, we measured the symmetric epipolar error w.r.t. the ground truth pose, translation and rotation errors, and the percentage of inliers. Fig. 3 in the main paper showed the rotation error and percentage of inliers, as observed for the St. Peter’s Square scene from the PhotoTourism dataset [23]. Here, Fig. 11 shows the same statistics, including translation and symmetric epipolar errors, for the St. Peter’s Square scene already used in the main paper (Fig. 11 (top row)), and two more scenes from the PhotoTourism dataset: Sacre Coeur (Fig. 11 (middle row)), and Temple Nara Japan (Fig. 11 (bottom row)).
@@ 7 Ablation studies
Validation of . We tested our -based solvers for different values of and measured their performance. In general, there is no common value of the shift that leads to the best results on all datasets. This is expected since the precision of the mean-point correspondence depends on many different factors, e.g., the viewing angles of the cameras, the type of the motion, the depth and spatial distributions of the 3D points, etc. We set the value for by evaluating their effects on the St. Peter’s Square scene from the PhotoTourism dataset [23], which we used for validation only and did not include it in the other results for PhotoTourism [23] in the paper. Tab. 3 shows how the different settings of the scale of the shift affect the accuracy of the -based solvers. Based on these experiments, we use as it typically provides the best or the second best results for all variants of the 4p3v(M)-based solvers. However, note that 4p3v(M)-based solvers achieve a very similar accuracy even with different settings of . Thus, we can conclude that the choice of is not critical. In some scenarios, the choice of the optimal parameter may be more scene-dependent and could potentially be set using learning-based approaches.
Refinement validation. We also perform validation of the total number of LM steps in the refinement (+R). Again, we used the St. Peter’s Square scene from the PhotoTourism dataset [23] for validation. The results of this experiment are shown in Fig. 12. We chose the value of 2 for other experiments as it provides good speed-accuracy trade-off across a range of RANSAC iterations. However, we note that other settings provide very similar performance.

### v5

@@ 5 Conclusion
We consider the highly challenging problem of relative pose estimation of three calibrated cameras from four correspondences. We propose novel approaches that solve the problem by utilizing solvers based on approximate geometry. The best performing solver uses a simple, yet novel strategy by using mean coordinates of three input points and points in their vicinity as an approximate correspondences. Extensive experiments show that our solvers achieve state-of-the-art performance on a large variety of real scenes. At the same time, our solvers are simple to implement, especially compared to the current state-of-the-art [Hruby_cvpr2022].
Acknowledgements. This work was funded by the Czech Science Foundation (GAČR) JUNIOR STAR Grant No. 22-23183M (supporting C.T., Y.D., and Z.K.), the Grant Agency of the Czech Technical University in Prague grant no. SGS23/173/OHK3/3T/13 (supporting C.T.), the EU NextGenerationEU through the Recovery and Resilience Plan for Slovakia under the project ”InnovAIte Slovakia, Illuminating Pathways for AI-Driven Breakthroughs” No. 09I02-03-V01-00029, the TERAIS project – a Horizon-Widera-2021 program of the European Union under the Grant agreement number 101079338 (supporting V.K. and Z.B.H.), and the EU Horizon 2020 project RICAIP grant agreement No. 857306 (supporting T.S.). Part of the research results was obtained using the computational resources procured in the national project National competence centre for high performance computing project code: 311070AKF2, funded by European Regional Development Fund, EU Structural Funds Informatization of society, Operational Program Integrated Infrastructure.
@@
This supplementary material provides additional details and experimental results promised in the main paper: Sec. 6 provides the proof on the bound of the epipolar error of the mean point correspondence mentioned in Sec. 3.1 of the main paper, additional synthetic and noise experiments that were discussed in Sec. 4 of the main paper, accuracy-speed trade-off results for two-view approximate geometry inside RANSAC (see Sec. 4 of the main paper), and additional details and plots on more scenes for the mean point correspondence accuracy (see Sec. 4 of the main paper). Sec. 7 contains ablation studies to validate our choices regarding the modifications discussed in Sec. 3.2 of the main paper. Sec. 8 provides results for the accuracy-speed trade-off for individual scenes from the PhotoTourism [IMC2020] and Cambridge Landmarks [kendall2015cambridge] datasets, evaluations of the three-view solvers for different thresholds inside RANSAC (see Sec. 4 of the main paper), runtimes of the proposed and state-of-the-art solvers for the 4p3v problem, semi-synthetic experiments for increasing outliers ratios on a scene from PhotoTourism [IMC2020], details and results on an alternative evaluation measure (see Sec. 4 of the main paper), and results with GC-RANSAC (see Sec. 4 of the main paper). 44footnotetext: ∗ Equal contribution
@@ 6 Approximate camera geometry
In this section, in addition to the experiments presented in Sec. 4 of the main paper (paragraph “Approximate camera geometry”), we present additional experiments and results to support our idea of estimating approximate geometry in the first two views. We start with the proof on the bound of the epipolar error of the mean-point correspondence used in the proposed 4p3v(M)-based solvers (see Sec 3.1 of the main paper). Next, Sec. 6.2 discusses why the mean-point correspondence provides an additional constraint (compared to the original point correspondences) that can be used to estimate the essential matrix. Then, to further assess the accuracy of the two-view variants of our approximate solvers (outside of RANSAC), i.e., 4p(A), 4p(M), and 4p(M), we design two additional synthetic experiments (see Sec. 6.3). The goal of these synthetic experiments is to study how the accuracy of approximate solutions varies with varying properties of the scene and the cameras. Moreover, Sec. 6.3 provides a synthetic noise experiment on data extracted from a real scene from the ETH3D dataset [Schops_2017_CVPR], outside of RANSAC for the three-view solvers. Lastly, Sec. 6.4 contains a speed-accuracy evaluation of the two-view variants of the proposed solvers, inside Poselib RANSAC, and in Sec. 6.5 we study the accuracy of the mean point correspondence (as in Fig. 3 of the main paper).
@@ 6.1 Proof of the bounds on the epipolar error
While the mean point correspondence used in the 4p3v(M)-based solvers can provide a good approximation of a correct correspondence, such a correspondence can be noisy. Note that all state-of-the-art 4p3v solvers (including 4p3v(HC) [Hruby_cvpr2022] and the solver from [DBLP:journals/ijcv/NisterS06]) rely on certain approximations without establishing theoretical proofs to quantify their accuracy. In the 4p3v(HC) solver [Hruby_cvpr2022], the failures that appear quite often are usually the results of tracking a geometrically incorrect solution inside the homotopy continuation method.555The solver is tracking only one from 272 solutions of the relaxed version of the 4p3v problem and this solution does not need to be a geometrically correct one. Thus, this solution can be arbitrarily far from the correct solution. The solver from [DBLP:journals/ijcv/NisterS06] requires sampling epipoles from a -degree curve on which the true epipole must lie. For any selected point on the curve of epipoles, the error of the sampled epipole is not bounded, since the true epipole can lie anywhere on the unbounded curve. The curve is unbounded since the epipole can be located arbitrarily far away from the image center based on the relative pose of the two cameras (up to infinity for sideways motion).
In contrast, the error of our mean point correspondence is bounded. The correspondence can be seen as a correspondence of points that are projections of the mean point of three 3D points. In this case, the error of both projections and can be computed as an error that is introduced by approximating the perspective projection using the para-perspective projection. The approximation error introduced by the para-perspective projection is studied in the literature and can be found e.g. in [Zhang2014, Horaud97].
@@ 6.2 Mean-point constraint
As already mentioned in the main paper, under the assumption of a para-perspective projection, i.e., of affine geometry, the mean point of three 3D points is projected to the mean points of the 3D points’ projections in both images [Zhang2014]. Thus, the mean point correspondence does not add a new constraint if used to estimate an affine camera. This can be easily shown. In the case of affine cameras, the essential matrix has the form
@@ 6.3 Synthetic Experiments
The error of the relative poses estimated with the proposed approximate 4p3v(M)-based and 4p3v(A)-based solvers depends on many aspects, e.g., the baseline and the view angles of the cameras w.r.t. the three points used to compute the mean point correspondence, the depth of these points, the size and shape of the triangles defined by the three points, the type of motion of the cameras, the depth of the scene and the distance of cameras from the scene, the level of noise in the correspondences, etc. Isolating the impact of the individual aspects, e.g., through experiments on synthetic data, is highly non-trivial (e.g., how to generate realistic synthetic scenarios that allow conclusions to generalize to real-world scenarios) and analysing the co-dependencies between different aspects on the overall performance seems to need a paper on its own. Moreover, the effect of approximation introduced by using para-perspective projection was already studied in the literature [Zhang2014, Horaud97].
@@
Additionally, we test the performance (outside of RANSAC) of our proposed approximate solvers and the state-of-the-art solvers for the 4p3v problem w.r.t. increasing image noise added to ground-truth correspondences extracted from a scene from the ETH3D dataset [Schops_2017_CVPR]. As an evaluation metric for the two-view geometry, we use the pose error measured as [IMC2020].
@@
Increasing distance of cameras to the 3D scene. It is known that the quality of the affine approximation, i.e., the approximation of the perspective projection using the para-perspective projection, depends on the distance of the points from the camera [Zhang2014]. Thus in the first experiment, we evaluate the performance of all solvers w.r.t. increasing distance of the cameras to the 3D scene.
@@
Noise experiments. We test the performance of our solvers and the state-of-the-art solvers w.r.t. increasing image noise. We used the SfM model of the botanical garden scene (randomly selected from all scenes) from the ETH3D dataset [Schops_2017_CVPR] to obtain instances of 5 points in three views by identifying images in the scene that share 3D points. Perfect noise-free image correspondences are generated by projecting the 3D points into the images. We then add increasing amounts of normally distributed noise to these correspondences. We generated more than 1k instances. Note that the 4p3v(HC) solver was trained on the ETH3D dataset [Schops_2017_CVPR].
@@
The results for increasing noise in the image points are shown in Fig. 9. The figure shows boxplots of the pose errors measured in the same way as in our experiments in the main paper (cf. Sec. 4 in the main paper), i.e., as .666Here is the error of the estimated relative rotation between cameras and , computed as the angle in the axis-angle representation of , and is the error of the estimated translation computed as the angle between the two unit vectors corresponding to the translations [IMC2020]. The errors are zoomed into an interesting interval and are shown as functions of varying Gaussian noise from px to px.
@@
Due to the approximations used in our proposed 4p3v(M)-based and 4p3v(A)-based solvers, these solvers exhibit non-zero errors for zero noise. However, at noise levels px, our 4p3v(M)+R solvers return comparable or even better (w/ ENM) results than the 5pt+P3P solver. For noise px, also the 4p3v(M)+R solver with ENM returns slightly more accurate poses than the 5pt+P3P solver with ENM. In general, the effect of increasing image noise is less visible for approximate 4p3v(M)-based and 4p3v(A)-based solvers. In this case, the error of the approximation is dominating the error introduced by the noise in the image correspondences. While for 4p3v(A)-based solvers the approximation error is dominant at all noise levels, for 4p3v(M)-based solvers, at noise px, the error introduced by the approximate mean point correspondence (and points in their vicinity in -based solvers) is suppressed by the error introduced by noise in the remaining point correspondences.777This can be seen from the comparable pose accuracy of the 5pt+P3P and 4p3v(M) solvers for px noise, and the comparable accuracy of the 5pt+P3P and 4p3v(M) solvers for px noise. Note that, although the pose errors for the 4p3V(A)+R+ENM solver are higher than those of the rest of the solvers, as shown in our real experiments, this solver still returns reasonably low errors to provide local optimization (LO) within RANSAC with a good initialization in real-world settings. Further, note that the 5pt+P3P solver samples one more point (real correspondence) in the first two cameras, and these points are affected only by the considered noise. This shows that the mean point correspondence used in the 4p3v(M)-based solver is a good approximation to a real correspondence. The recent state-of-the-art 4p3v(HC) solver [Hruby_cvpr2022] is failing in about 50% of the instances for noiseless data, even though the solver was trained on the ETH3D dataset. Thus, the median errors are significantly larger than the median errors of the remaining solvers.
@@ 6.4 Two-view approximate solutions inside RANSAC
Fig. 10 shows the speed-accuracy evaluation of the solvers for the problem of two view relative pose estimation evaluated using PoseLib RANSAC [PoseLib] on 12 scenes from the Phototourism dataset [IMC2020] (excluding St. Peter’s Square) with pairwise point correspondences obtained using [detone2018superpoint, lindenberger2023lightglue]. The statistic reported is the AUC@10∘ of the pose error for a varied number of RANSAC iterations () and a fixed epipolar threshold of 5px. The upper figure is a zoom-in of the lower figure to an interesting interval, where differences between the solvers are more visible. Although all proposed solvers (except the 4p(A) solver without ENM) have a performance comparable to that of the state-of-the-art two view 5pt solver, the 5pt solver is the best performing one for the two-view scenario. This result is not surprising, given the well-known good performance (in terms of speed and accuracy) of the 5pt solver and not very high outlier contamination of the data (for which sampling one point less would have potentially had a more visible effect). It also indicates that the proposed modifications, i.e., the filtering +F the and refinement +R, for the three-view scenario are important and are making the proposed 4p3v approximate solvers practical and more precise than the 5pt+P3P sovler.
@@ 6.5 Accuracy of the mean point correspondence
Fig. 3 in the main paper showed results obtained by establishing correspondences between the mean of the triangle in one image and various points in the triangle in the second image. We expressed points in the second triangle via their barycentric coordinates and uniformly sample barycentric coordinates , such that (ensuring points inside the triangle). The 3rd coordinate is given as . For each correspondence, we measured the symmetric epipolar error w.r.t. the ground truth pose, translation and rotation errors, and the percentage of inliers. Fig. 3 in the main paper showed the rotation error and percentage of inliers, as observed for the St. Peter’s Square scene from the PhotoTourism dataset [IMC2020]. Here, Fig. 11 shows the same statistics, including translation and symmetric epipolar errors, for the St. Peter’s Square scene already used in the main paper (Fig. 11 (top row)), and two more scenes from the PhotoTourism dataset: Sacre Coeur (Fig. 11 (middle row)), and Temple Nara Japan (Fig. 11 (bottom row)).
@@ 7 Ablation studies
Validation of . We tested our -based solvers for different values of and measured their performance. In general, there is no common value of the shift that leads to the best results on all datasets. This is expected since the precision of the mean-point correspondence depends on many different factors, e.g., the viewing angles of the cameras, the type of the motion, the depth and spatial distributions of the 3D points, etc. We set the value for by evaluating their effects on the St. Peter’s Square scene from the PhotoTourism dataset [IMC2020], which we used for validation only and did not include it in the other results for PhotoTourism [IMC2020] in the paper. Tab. 4 shows how the different settings of the scale of the shift affect the accuracy of the -based solvers. Based on these experiments, we use as it typically provides the best or the second best results for all variants of the 4p3v(M)-based solvers. However, note that 4p3v(M)-based solvers achieve a very similar accuracy even with different settings of . Thus, we can conclude that the choice of is not critical. In some scenarios, the choice of the optimal parameter may be more scene-dependent and could potentially be set using learning-based approaches.
Refinement validation. We also perform validation of the total number of LM steps in the refinement (+R). Again, we used the St. Peter’s Square scene from the PhotoTourism dataset [IMC2020] for validation. The results of this experiment are shown in Fig. 12. We chose the value of 2 for other experiments as it provides good speed-accuracy trade-off across a range of RANSAC iterations. However, we note that other settings provide very similar performance.